The bumper cars A and B in Fig. Q5 below each have a mass of 150 kg and are coasting with the velocities shown before they freely collide head on. If no energy is lost during the collision, determine their velocities after collision. [10]


1
Expert's answer
2021-10-08T09:23:57-0400

"m_a=m_b = 150 \; kg \\

(v_A)_1=3 \;m/s \\

(v_B)_1 = 2 \;m/s"

No energy lost

"E_l=0 \\

e=1 \\

e = \frac{-[(V_B)_2 -(V_A)_2]}{(V_B)_1 -(V_A)_1}"

A and B are in opposite direction

"(V_A)_1 = 3 \;m/s \\

(V_B)_1 = -2\;m/s \\

1 = \frac{-[(V_B)_2 -(V_A)_2]}{-2-3} \\

(V_B)_2 -(V_A)_2 = 5"

Conservation of momentum

"m_a(V_A)_1 +m_b(V_B)_1 =m_a(V_A)_2 +m_b(V_B)_2 \\

m_a=m_b \\

(V_A)_1 -(V_B)_1 = (V_A)_2 +(V_B)_2 \\

3-2 = (V_A)_2 +(V_B)_2 \\

1 = (V_A)_2 +(V_B)_2 \\

5 = (V_B)_2 -(V_A)_2 \\

(V_A)_2 = -2 \;m/s \\

(V_B)_2 = 3 \;m/s"